OpenStudy (anonymous):

Simplify ?! (2+√10)(2-√10)

OpenStudy (anonymous):

@thomaster could you help?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jim_thompson5910 could you help

OpenStudy (accessdenied):

Have you tried anything here so far? Or had any ideas that might work out but you' re unsure of?

OpenStudy (anonymous):

@AccessDenied Not really. because 10 isn't a perfect square

OpenStudy (accessdenied):

When I look at this, I think about a factoring trick: difference of squares? is that one familiar? \(a^ 2 - b^ 2 = (a + b)(a -b) \)

OpenStudy (anonymous):

@AccessDenied yeah i think so..

OpenStudy (accessdenied):

So the reason I bring that up is, the right side of that equation matches what we have now if we called a = 2 and b = sqrt(10). And on the left side, we end up squaring the 2 and that pesky square root of 10 as well. Or filling in the details: \( 2^ 2 - \sqrt{10}^2 = (2 + \sqrt{10})(2 - \sqrt{10}) \) Can you see how to simplify the left side now? (Square root and square cancel first)

OpenStudy (anonymous):

@AccessDenied so then we just cancel out the squaring of 2 and 10 so it would just be 2-√10?

OpenStudy (anonymous):

@AccessDenied hello?

OpenStudy (accessdenied):

You would not change 2^2. The square root on the 10 and the square cancel out each other as they are inverses. Like this: \( 2^ 2 - \sqrt{10}^2 = 2^2 - 10 \)

OpenStudy (anonymous):

@AccessDenied oh okay. so then we would end up with 2^2 - 10?

OpenStudy (accessdenied):

Well, you can simplify more from there. Order of operations, we do the exponents first: 2^ 2 = 2*2 = 4. Then subtract 4 - 10. Makes sense?

OpenStudy (anonymous):

@AccessDenied yes okay. so then the answer would be -6.?

OpenStudy (accessdenied):

Yes. :)

OpenStudy (anonymous):

@AccessDenied yay thank you for your help.:)

OpenStudy (accessdenied):

You' re welcome!